(a) A rectangular plot is to be enclosed by fencing wire. Assume that 56 𝑚 of fencing
wire are available.
(i) Show that the area of the plot, in 𝑚^2
, is given by the expression
𝐴 = −𝑥^2 + 28𝑥,
where 𝑥 is the length of the plot.
If someone could explain how I am to work these out that would be great i've been stuck on this for hours and can't find anything

Answer by Alan3354(69443)

(a) A rectangular plot is to be enclosed by fencing wire. Assume that 56 meters of fencing wire are available.
(i) Show that the area of the plot, in m^2
is given by the expression
Area = -x^2 + 28x,
where is the length of the plot.
-----
If the 56 meters is to be 4 sides or a rectangle:
56 = perimeter = 2W + 2L
---
x = width
28 - x = length
------------
Area = W*L = x*(28-x)
= -x^2 + 28x